# Optimistic Safety for Linearly-Constrained Online Convex Optimization

### v1

@@ Abstract
The setting of online convex optimization (OCO) under unknown constraints has garnered significant attention in recent years. In this work, we consider a version of this problem with static linear constraints that the player receives noisy feedback of and must always satisfy. By leveraging our novel design paradigm of optimistic safety, we give an algorithm for this problem that enjoys $\tilde{\mathcal{O}}(\sqrt{T})$ regret. This improves on the previous best regret bound of $\tilde{\mathcal{O}}(T^{2/3})$ while using only slightly stronger assumptions of independent noise and an oblivious adversary. Then, by recasting this problem as OCO under time-varying stochastic linear constraints, we show that our algorithm enjoys the same regret guarantees in such a setting and never violates the constraints in expectation. This contributes to the literature on OCO under time-varying stochastic constraints, where the state-of-the-art algorithms enjoy $\tilde{\mathcal{O}}(\sqrt{T})$ regret and $\tilde{\mathcal{O}}(\sqrt{T})$ violation when the constraints are convex and the player receives full feedback. Additionally, we provide a version of our algorithm that is more computationally efficient and give numerical experiments comparing it with benchmark algorithms.
@@ 1 Introduction
Online convex optimization (OCO), formalized by Zinkevich (2003), is a setting where, in each round, a player chooses a vector action, suffers the cost of that action according to an adversarily-chosen cost function, and subsequently observes the cost function. This setting has found broad applicability in various fields including online advertising McMahan et al. (2013), network routing Awerbuch & Kleinberg (2008), portfolio management Cover (1991) and resource allocation Yu & Neely (2019). However, despite its significance, this fundamental setting assumes that any constraints on the player’s actions are known a priori, which is not the case in many real-world applications. As a result, the problem of OCO under unknown constraints has received considerable attention in the last decade, e.g. Mannor et al. (2009); Yu et al. (2017); Cao & Liu (2018); Guo et al. (2022).
Two distinct directions have emerged in the literature on OCO under unknown constraints. The first direction considers time-varying constraints that are revealed after the player commits to an action, while the second direction considers unknown static constraints that the player receives incomplete feedback on. When the constraints are time-varying, the player cannot hope to always satisfy the constraint so it is instead natural to bound the cumulative constraint violation [^1] Here, we use $g_{t}$ to refer to the constraint in the $t$th round. $\sum_{t=1}^{T}g_{t}(x_{t})$ that the player incurs. However, it has been shown that sublinear regret and sublinear violation cannot be simultaneously achieved Mannor et al. (2009), so several works have instead focused on the case with iid constraints and provided algorithms with $\tilde{\mathcal{O}}(\sqrt{T})$ regret and $\tilde{\mathcal{O}}(\sqrt{T})$ violation Yu et al. (2017); Wei et al. (2020); Castiglioni et al. (2022). On the other hand, the setting with OCO under fixed unknown constraints enjoys $\tilde{\mathcal{O}}(T^{2/3})$ regret and zero instantaneous constraint violation (i.e. $g(x_{t})\leq 0\ \forall t$) when the constraints are linear and the player receives noisy constraint feedback at the chosen actions Chaudhary & Kalathil (2022); Chang et al. (2023).
Ref. Const. Var. Const. Type Const. Feedback Regret Cumul. Viol. Instant. Viol.
[1],[2] static linear noisy bandit $\tilde{\mathcal{O}}(T^{2/3})$ 00
Thm. 4.1 static linear noisy bandit $\tilde{\mathcal{O}}(\sqrt{T})$ 00
[3],[4],[5] stochastic convex full $\tilde{\mathcal{O}}(\sqrt{T})$ $\tilde{\mathcal{O}}(\sqrt{T})$
Prop. 5.1 stochastic linear bandit $\tilde{\mathcal{O}}(\sqrt{T})$ 00
Table 1: Existing work and novel contributions on OCO under unknown static constraints and OCO under stochastic constraints. Our contributions are highlighted in gray. Prior work is referenced as: [1] = Chaudhary & Kalathil (2022), [2] = Chang et al. (2023), [3] = Yu et al. (2017), [4] = Wei et al. (2020), [5] = Castiglioni et al. (2022). In the “instantaneous violation” column, the stated guarantees hold with high probability for static constraints and in expectation for stochastic constraints.
We contribute to this literature as summarized in Table 1 and in the following:
  • We give the first algorithm with $\tilde{\mathcal{O}}(\sqrt{T})$ regret for OCO under unknown linear constraints with noisy feedback. With only slightly stronger assumptions of independent noise and an oblivious adversary, this improves on the best existing $\tilde{\mathcal{O}}(T^{2/3})$ regret attained by Chaudhary & Kalathil (2022) and Chang et al. (2023).
  • We show that, with slight modification, our algorithm enjoys $\tilde{\mathcal{O}}(\sqrt{T})$ regret and zero instantaneous violation under stochastic linear constraints and bandit constraint feedback. To put this in context, the state-of-the-art guarantees for OCO under stochastic constraints are $\tilde{\mathcal{O}}(\sqrt{T})$ regret and $\tilde{\mathcal{O}}(\sqrt{T})$ cumulative violation Yu et al. (2017); Castiglioni et al. (2022) when the constraints are convex and the player recieves full feedback.
These contributions are thanks to our novel approach of using optimistic estimates of the constraints (outer-approximations of the constraint set) to identify actions that are low-regret, but potentially unsafe, and then using pessimistic estimates of the constraints (inner-approximations of the constraint set) to find the scalings on these actions that will ensure constraint satisfaction. Ultimately, this results in decision-making that is simultaneously low-regret and safe. We call this design paradigm optimistic safety and foresee it being broadly applicable to learning and optimization under uncertain constraints.
@@ 1.1 Overview
In Section 2, we define relevant notation and then give the problem setup and assumptions. Then, we design a meta-algorithm in Section 3 that leverages the idea of optimistic safety by using a generic online optimization algorithm to choose actions from the optimistic action set and then scales these actions down into the pessimistic set before playing them. In Section 4, we show that this meta-algorithm ensures $\tilde{\mathcal{O}}(d\sqrt{T})$ regret and zero constraint violation. We then recast the problem as OCO under stochastic constraints in Section 5, and show that our proposed approach enjoys $\tilde{\mathcal{O}}(d\sqrt{T})$ regret guarantees and zero expected constraint violation. In Section 6, we address computational inefficiencies in our original algorithm by giving a version that only requires $2d$ convex projections in each round and enjoys $\tilde{\mathcal{O}}(d^{3/2}\sqrt{T})$ regret. Lastly, we give simulation results in Section 7 that verify the theoretical claims and show improved performance compared to baseline approaches.
@@ 1.2 Related work
This work is related to works on OCO with unknown linear constraints, OCO with time-varying constraints, projection-free OCO, safe learning and online control. We provide separate discussion for each of these fields in the following.
OCO with Unknown Constraints
The problem of OCO under unknown linear constraints and noisy feedback was first studied by Chaudhary & Kalathil (2022) who gave an algorithm with $\tilde{\mathcal{O}}(T^{2/3})$ regret guarantees. In comparison to Chaudhary & Kalathil (2022), we use the slightly stronger assumptions of independent noise and oblivious adversary as discussed in Section 2.3. The approach taken by Chaudhary & Kalathil (2022) was then extended to the distributed and nonconvex setting in Chang et al. (2023), where an $\tilde{\mathcal{O}}(T^{2/3}+T^{1/3}C_{T})$ regret bound was given with $C_{T}$ as the path length of the best minimizer sequence.
@@
For OCO with time-varying constraints, Mannor et al. (2009) showed that no algorithm can simultaneously achieve both sublinear constraint violation and sublinear regret against the best action that satisfies the constraints on average. Following this negative result, considerable effort has been focused on ways to weaken the setting to guarantee sublinear regret and violation, e.g. Neely & Yu (2017); Sun et al. (2017); Liakopoulos et al. (2019); Cao & Liu (2018); Yi et al. (2022); Guo et al. (2022); Castiglioni et al. (2022). Most relevantly, several works study the setting in which the constraints are bounded convex functions that are sampled iid and guarantee $\tilde{\mathcal{O}}(\sqrt{T})$ regret and cumulative violation Yu et al. (2017); Wei et al. (2020); Castiglioni et al. (2022). In Section 5, we guarantee $\tilde{\mathcal{O}}(\sqrt{T})$ regret and zero expected instantaneous violation in a related setting where the constraints are linear with subgaussian tails and the player receives bandit feedback of the constraint. Note that our assumptions of subgaussianity and bandit feedback are weaker than bounded and full feedback as considered by Yu et al. (2017), Wei et al. (2020) and Castiglioni et al. (2022), although our assumption of linearity is stronger than their assumption of convexity.
Projection-free OCO
The problem of OCO under unknown constraints has similar goals to that of projection-free OCO, which is focused on methods for OCO that do not require the costly projection operation required for classical methods, such as OGD Zinkevich (2003). A large number of projection-free OCO methods use the cheaper linear optimization oracle instead of the projection oracle and often use variants of the Franke-Wolfe method, e.g. Hazan & Kale (2012); Garber & Hazan (2016); Hazan & Minasyan (2020); Kretzu & Garber (2021). Another line of work avoids projections by allowing some constraint violation and uses similar approaches to OCO with time-varying constraints, e.g. Mahdavi et al. (2012); Yu & Neely (2020); Guo et al. (2022). Some other approaches use the membership oracle Mhammedi (2022) or the constraint function and gradient at any point Levy & Krause (2019). The literature on projection-free OCO is distinct from OCO under unknown constraints in that they either allow constraint violation or consider different types of oracles that use the constraints, e.g. membership oracle, constraint function value and gradient at any point, or linear optimization oracle.
@@
Our setting falls under the umbrella of “safe learning,” which often deals with learning problems that have uncertain constraints. Some examples of such problems are constrained MDPs Achiam et al. (2017); Wachi & Sui (2020); Liu et al. (2021); Amani et al. (2021), safe Gaussian process bandits Sui et al. (2015, 2018), safe optimization with uncertain constraints Usmanova et al. (2019); Fereydounian et al. (2020) and safe linear bandits Amani et al. (2019); Moradipari et al. (2021); Pacchiano et al. (2021). The most relevant to our setting is safe linear bandits Moradipari et al. (2021); Pacchiano et al. (2021) as the two settings have the same constraint feedback and both concern regret minimization. However, the safe linear bandit is distinct from the setting we consider as it has stochastic linear cost functions and bandit feedback versus the adversarial convex cost functions and full cost feedback used in our setting. Also, note that the setting we study in the present work generalizes the safe optimization setting in Usmanova et al. (2019) as they consider the same constraints and constraint feedback with a static cost function.
Online Control
2 Preliminaries
@@ 2.1 Notation
2.2 Problem Setup
We study an online convex optimization problem with an unknown linear constraint. This problem is played over $T$ rounds, where in each round $t\in[T]$, a player chooses an action $x_{t}$ from the convex action set $\mathcal{X}\subseteq\mathbb{R}^{d}$ and, subsequently, suffers the cost $f_{t}(x_{t})$ according to an adversarily-chosen convex function $f_{t}:\mathcal{X}\rightarrow\mathbb{R}$. The player then observes the cost function $f_{t}$ and receives noisy feedback on the constraint $Ax_{t}+\epsilon_{t}$ where $A\in\mathbb{R}^{n\times d}$ is the constraint matrix and $\epsilon_{t}$ is a noise term.
The player aims to ensure that the constraint $Ax_{t}\leq b$ is always satisfied despite the fact that the constraint matrix $A$ is unknown. At the same time, the player also aims to minimize her regret with respect to the best constraint-satisfying action in hindsight, i.e.
@@ 2.2 Problem Setup
2.3 Assumptions
Our approach to this setting relies on the following assumptions. We first assume that the cost functions have bounded gradients and the action set is compact, which are standard in OCO (e.g. Zinkevich (2003); Hazan et al. (2016))
Assumption 2.1 (Bounded gradients).
@@ 2.3 Assumptions
Assumption 2.2 (Compact action set).
@@
We then assume that the noise terms are independent and subgaussian, the constraint matrix is bounded and that there is an initial safe action that is known. These match the assumptions used in Chaudhary & Kalathil (2022); Chang et al. (2023), except that we use the mild additional assumption that the noise terms are independent.
Assumption 2.3 (Subgaussian noise).
The noise $(\epsilon_{t})_{t\in[T]}$ is a sequence of independent random variables that are element-wise $\rho$-subgaussian, i.e. it holds for all $t\in[T]$ and $i\in[n]$ that $\mathbb{E}[\epsilon_{t,i}]=0$ and $\mathbb{E}[\exp(\lambda\epsilon_{t,i})]\leq\exp(\frac{\lambda^{2}\rho^{2}}{2}),\forall\lambda\in\mathbb{R}$ where $\epsilon_{t,i}$ is the $i$th element of $\epsilon_{t}$.
Assumption 2.4 (Bounded constraint).
@@
Assumption 2.5 (Initial safe action).
[…]
Lastly, we assume that the adversary is oblivious such that the cost functions are chosen independently of the player’s actions and the noise terms.
Assumption 2.6 (Oblivious adversary).
The adversary is oblivious and therefore the cost functions are fixed in advance.

### v3

# Optimistic Safety for Online Convex Optimization with Unknown Linear Constraints

@@ Abstract
We study the problem of online convex optimization (OCO) under unknown linear constraints that are either static, or stochastically time-varying. For this problem, we introduce an algorithm that we term Optimistically Safe OCO (OSOCO) and show that it enjoys $\tilde{\mathcal{O}}(\sqrt{T})$ regret and no constraint violation. In the case of static linear constraints, this improves on the previous best known $\tilde{\mathcal{O}}(T^{2/3})$ regret under the same assumptions. In the case of stochastic time-varying constraints, our work supplements existing results that show $\mathcal{O}(\sqrt{T})$ regret and $\mathcal{O}(\sqrt{T})$ cumulative violation under more general convex constraints and a different set of assumptions. In addition to our theoretical guarantees, we also give numerical results that further validate the effectiveness of our approach.
@@ 1 Introduction
Online convex optimization (OCO), formalized by Zinkevich (2003), is a setting where, in each round $t\in[T]$, a player chooses a vector action $x_{t}$, suffers the cost of that action according to an adversarially-chosen cost function $f_{t}(x_{t})$, and subsequently observes the cost function $f_{t}$. This setting has found broad applicability in various fields including online advertising (McMahan et al., 2013), network routing (Awerbuch and Kleinberg, 2008), portfolio management (Cover, 1991) and resource allocation (Yu and Neely, 2019). However, despite its significance, this fundamental setting assumes that any constraints on the player’s actions are known a priori, which is not the case in many real-world applications. As a result, the problem of OCO under unknown constraints has received considerable attention in the last decade, e.g. (Mannor et al., 2009; Yu et al., 2017; Cao and Liu, 2018; Guo et al., 2022).
Within the broader literature on OCO with unknown constraints, this paper focuses on two important research directions: (1) OCO with stochastic constraints, where there are different constraints in each round $g_{t}(x_{t})\leq 0\ \forall t$ that are independent and identically-distributed (i.i.d.) and the player receives feedback on $g_{t}$ after choosing each action $x_{t}$, and (2) OCO with static constraints, where there is a static and deterministic constraint $g(x_{t})\leq 0\ \forall t$ for which the player receives noisy feedback $g(x_{t})+\epsilon_{t}$ after choosing each action $x_{t}$. In OCO with stochastic constraints, existing works typically leverage primal-dual algorithms to ensure both sublinear regret and sublinear cumulative violation $\sum_{t=1}^{T}g_{t}(x_{t})$. These algorithms operate by jointly iterating over both the actions (primal variables) and constraint penalties (dual variables), and have been shown to ensure $\mathcal{O}(\sqrt{T})$ [^1] For ease of presentation in the introduction, we use $\mathcal{O}(\cdot)$ to hide non-$T$ quantities, and $\tilde{\mathcal{O}}(\cdot)$ to further hide $log$ factors. expected regret and $\mathcal{O}(\sqrt{T})$ expected cumulative violation (Yu et al., 2017; Wei et al., 2020; Castiglioni et al., 2022; Yu et al., 2023). On the other hand, the works on OCO with static constraints ensure no violation $g(x_{t})\leq 0\ \forall t$ with high probability, by explicitly estimating the constraints and then only choosing actions that are verifiably-safe (Chaudhary and Kalathil, 2022; Chang et al., 2023). However, the approach in Chaudhary and Kalathil (2022) and Chang et al. (2023) requires a pure-exploration phase of $\Theta(T^{2/3})$ rounds and therefore incurs a less favorable regret of $\tilde{\mathcal{O}}(T^{2/3})$.
In this work, we introduce an algorithm for both static and stochastic constraints that directly estimates the constraints, and therefore ensures no violation, while avoiding the costly exploration phase used in prior works (Chaudhary and Kalathil, 2022; Chang et al., 2023). Precisely, our algorithm uses optimistic estimates of the constraints (outer-approximations of the feasible set) to identify actions that are low-regret, but potentially unsafe, and then pessimistic estimates of the constraints (inner-approximations of the feasible set) to find the scalings on these actions that will ensure constraint satisfaction. This approach, which we call optimistic safety, allows for the constraint to be learned simultaneous to the regret minimization and therefore does not require a pure-exploration phase as in prior works. Ultimately, our approach yields the following results:
  • In OCO with static linear constraints and noisy constraint feedback, we show that our algorithm enjoys $\tilde{\mathcal{O}}(\sqrt{T})$ regret and no violation, in either high probability (Theorem 1) or in expectation (Proposition 1). Our high probability guarantees improve on the $\tilde{\mathcal{O}}(T^{2/3})$ regret attained by prior work on this problem, under the same assumptions and violation guarantees (Chaudhary and Kalathil, 2022; Chang et al., 2023).
  • We then extend our results to OCO with stochastic linear constraints $g_{t}$, where the player receives bandit feedback on the constraint $g_{t}(x_{t})$ (Section 6). We show $\tilde{\mathcal{O}}(\sqrt{T})$ expected regret and no violation in expectation $\mathbb{E}[g_{t}(x_{t})]\leq 0\ \forall t$ in this setting (Corollary 1). This supplements existing results on OCO with stochastic convex constraints that show $\mathcal{O}(\sqrt{T})$ expected regret and $\mathcal{O}(\sqrt{T})$ expected cumulative violation $\mathbb{E}[\sum_{t}g_{t}(x_{t})]$, although with a different set of assumptions (see Section 6.1) (Yu et al., 2017; Wei et al., 2020; Castiglioni et al., 2022; Yu et al., 2023).
@@ 1.1 Overview
In Section 2, we give the problem setup and assumptions. Then, we design a meta-algorithm in Section 1 that leverages the idea of optimistic safety. In Section 4, we show that this meta-algorithm ensures $\tilde{\mathcal{O}}(d\sqrt{T})$ regret and no constraint violation in either high-probability or expectation. In Section 5, we improve the computational efficiency of our original algorithm by giving a version that only requires $2d$ convex projections in each round, although with a slightly larger dependence on the dimension in the regret bound, i.e. $\tilde{\mathcal{O}}(d^{3/2}\sqrt{T})$. In Section 6, we extend our approach to the setting of OCO with stochastic linear constraints and bandit constraint feedback and show matching regret bounds in expectation and no expected violation. Lastly, we give simulation results in Section 7 that showcase the empirical performance of our algorithm.
@@ 1.2 Related work
In the following, we discuss prior work on OCO with unknown static constraints, OCO with time-varying constraints, and safe learning. Lastly, we discuss a concurrent work. In Appendix A, we provide additional discussion on OCO with long-term constraints, projection-free OCO, and online control.
OCO with Unknown Static Constraints
The problem of OCO with unknown linear constraints and noisy constraint feedback was first studied by Chaudhary and Kalathil (2022) who gave an algorithm with $\tilde{\mathcal{O}}(dT^{2/3})$ regret and no violation $g(x_{t})\leq 0\ \forall t$, with high probability. In the same setting, we give an algorithm that enjoys $\tilde{\mathcal{O}}(d\sqrt{T})$ regret and no violations with high probability, and a more efficient variant that enjoys $\tilde{\mathcal{O}}(d^{3/2}\sqrt{T})$ regret. With a different choice of algorithm parameters, we also show matching bounds on the expected regret and show no expected violation $\mathbb{E}[g_{t}(x_{t})]\leq 0$. The approach taken by Chaudhary and Kalathil (2022) was then extended to the distributed and nonconvex setting in Chang et al. (2023), where an $\tilde{\mathcal{O}}(dT^{2/3}+T^{1/3}C_{T})$ high-probability regret bound was given with $C_{T}$ as the path length of the best minimizer sequence. Another variant of OCO with unknown static constraints was considered by Hutchinson and Alizadeh (2024), although they assume that the constraint function is strongly-convex, and thus their setting is distinct from the one we consider here.
@@
For OCO with time-varying constraints that are chosen adversarially, Mannor et al. (2009) showed that no algorithm can simultaneously achieve both sublinear cumulative violation and sublinear regret against the best action that satisfies the constraints on average. Following this negative result, considerable effort has been focused on ways to weaken the setting to guarantee sublinear regret and cumulative violation, e.g. (Neely and Yu, 2017; Sun et al., 2017; Chen et al., 2017; Liakopoulos et al., 2019; Cao and Liu, 2018; Yi et al., 2022; Guo et al., 2022; Castiglioni et al., 2022; Kolev et al., 2023). Most relevant to this paper, several works study the setting in which the constraints are convex functions that are stochastic and i.i.d., and give algorithms with $\mathcal{O}(\sqrt{T})$ expected regret and $\mathcal{O}(\sqrt{T})$ expected cumulative violation when the player receives full feedback (Yu et al., 2017; Wei et al., 2020; Castiglioni et al., 2022) and $\mathcal{O}(d\sqrt{T})$ expected regret and violation when the player receives two-point bandit feedback (Yu et al., 2023). In the present work, we supplement this literature by consider a setting (in Section 6) where the constraints are i.i.d. linear functions and the player receives one-point bandit feedback on the constraint (but full feedback on the cost). In this setting, we guarantee $\tilde{\mathcal{O}}(d\sqrt{T})$ expected regret (or $\tilde{\mathcal{O}}(d^{3/2}\sqrt{T})$ for our more efficient version) and no expected violation. We additionally assume that the player knows an action that is strictly feasible in expectation and its expected constraint value, as discussed in Section 6.1.
@@
Our setting falls under the classification of “safe learning,” which often deals with learning problems that have uncertain constraints. Some examples of such problems are constrained MDPs e.g. (Achiam et al., 2017; Wachi and Sui, 2020; Liu et al., 2021; Amani et al., 2021; Ghosh et al., 2024), safe Gaussian process bandits e.g. (Sui et al., 2015, 2018; Losalka and Scarlett, 2024), safe optimization with uncertain constraints e.g. (Usmanova et al., 2019; Fereydounian et al., 2020) and safe stochastic linear bandits e.g. (Amani et al., 2019; Moradipari et al., 2021; Pacchiano et al., 2021; Hutchinson et al., 2024). The most relevant to the present work is the problem of safe stochastic linear bandits (Moradipari et al., 2021; Pacchiano et al., 2021; Hutchinson et al., 2024) as it has the same constraint formulation as OCO with static linear constraints and both concern regret minimization. However, the safe stochastic linear bandit setting has stochastic linear costs with bandit feedback, while OCO with static linear constraints has adversarial convex costs with full cost feedback. Note that our approach draws inspiration from the safe stochastic linear bandit work Hutchinson et al. (2024), although we emphasize that OCO with static linear constraints has fundamental challenges that are not present in the safe stochastic linear bandit problem. Most notably, OCO with static linear constraints has the challenges posed by time-varying and nonconvex action sets in OCO, for which we use a phased-based approach and a novel OCO algorithm for nonconvex action sets.
@@ 1.3 Notation
2 Problem Setup
The problem is played over $T$ rounds, where in each round $t\in[T]$, a player chooses an action $x_{t}$ from the convex action set $\mathcal{X}\subseteq\mathbb{R}^{d}$ and, simultaneously, an adversary chooses the convex cost function $f_{t}:\mathcal{X}\rightarrow\mathbb{R}$ according to the actions chosen by the player in previous rounds. The player then suffers the cost $f_{t}(x_{t})$, observes the cost function $f_{t}$ and receives noisy constraint feedback $y_{t}:=Ax_{t}+\epsilon_{t}$. The matrix $A\in\mathbb{R}^{n\times d}$ is the unknown and deterministic constraint matrix, and the vector $\epsilon_{t}\in\mathbb{R}^{n}$ is a zero-mean random noise. Additionally, there is a deterministic constraint limit $b\in\mathbb{R}^{n}$ that is known to the player. [^2] The assumption that $b$ is known is also used in prior work on this problem (Chaudhary and Kalathil, 2022; Chang et al., 2023).
We aim either to ensure that there is no constraint violation in high-probability or no constraint violation in expectation. These two types of guarantees are summarized as follows:
  • High-probability: $Ax_{t}\leq b$ for all $t\in[T]$, with high probability.
  • Expectation: $\mathbb{E}[Ax_{t}]\leq b$ for all $t\in[T]$.
In either case, we additionally aim to minimize the regret with respect to the best constraint-satisfying action in hindsight, i.e.
@@ 2 Problem Setup
Our approach to this setting relies on the following assumptions. We first assume that the costs have bounded gradients and the action set is compact, which are standard in OCO (e.g. Zinkevich (2003); Hazan et al. (2016)).
Assumption 1 (Bounded gradients).
@@
Assumption 2 (Compact action set).
@@
We further assume that elements of $b$ are positive, the noise is conditionally sub-gaussian, and that the constraint parameter $A$ is bounded. These match the assumptions used in Chaudhary and Kalathil (2022) and Chang et al. (2023).
Assumption 3 (Initial safe action).
It holds that $\mathbf{0}\in\mathcal{X}$ and $b_{\mathrm{min}}:=\min_{i\in[n]}b_{i}>0$ where $b_{i}$ is the $i$th element of $b$ [^3] If a nonzero safe action $x^{s}$ and its constraint value $Ax^{s}$ are given (as in Chaudhary and Kalathil (2022)), then the problem can be shifted such that the safe action is at the origin. This will not impact the other problem parameters, except that $D$ may be increased by a factor $\leq 2$.
Assumption 4 (Conditionally-subgaussian noise).
The noise terms $(\epsilon_{t})_{t\in[T]}$ are element-wise conditionally $\rho$-subgaussian, i.e. it holds for all $t\in[T]$ and $i\in[n]$ that $\mathbb{E}[\epsilon_{t,i}|x_{1},\epsilon_{1},...,\epsilon_{t-1},x_{t}]=0$ and $\mathbb{E}[\exp(\lambda\epsilon_{t,i})|x_{1},\epsilon_{1},...,\epsilon_{t-1},x_{t}]\leq\exp(\frac{\lambda^{2}\rho^{2}}{2}),\forall\lambda\in\mathbb{R}$ where $\epsilon_{t,i}$ is the $i$th element of $\epsilon_{t}$.
Assumption 5 (Bounded constraint).
@@
3 Algorithm